OpenStudy (anonymous):

how do you solve this using log?: 10000 = 2 x 3 ^(n-1)

OpenStudy (anonymous):

log 10000=1og[2*3^(n-1)] log 10000=log2+(n-1)log3 log 10000-log 2=(n-1)log3 (log 10000-log 2)/log 3=n-1 7.75=n-1 8.75=n